OpenStudy (anonymous):

Decide whether the sequence is arithmetic,geometric , or neither. 3,12,21,30,...

OpenStudy (anonymous):

If the pattern uses multiplication and division then it is geometric. If the pattern uses addition or subtraction then it is arithmetic

OpenStudy (anonymous):

for arithmetic , Tn = a + (n-1)d , a= 1st term = 3 , d= 2nd -1st term = 9 T2 = 3 + 9 = 12 T3 = 3 + 2 (9) =3 + 18 = 21 T4 = 3 + 3(9) = 3 + 27 = 30 so Yes the sequence is arithmetic... :)

OpenStudy (anonymous):

you see that it keeps gaining 9, so it's adding, therefore arithmetic (:

OpenStudy (anonymous):

oh wow thanks babes! both so helpful:)

OpenStudy (anonymous):

no problem doll! (:

OpenStudy (anonymous):

no problem :)